# Phase transitions from linear to nonlinear information processing in neural networks

### v1

@@ IV.4 Other tasks
To demonstrate the robustness of the transition discussed above, we evaluate the information processing capacity using alternative benchmark tasks. One such task is the widely used nonlinear autoregressive moving average (NARMA) task Rodan-11 ; Appeltant-11 , defined by
$z(t+1)=a_{1}z(t)+a_{2}z(t)\sum_{i=0}^{m-1}z(t-i)+a_{3}u(t-m+1)u(t)+a_{4},$ (20)
with parameters $m=10$, $a_{1}=0.3$, $a_{2}=0.05$, $a_{3}=0.2$, and $a_{4}=0.05$. We also consider the delay task, defined by $z(t)=u(t-k)$, which simply requires the network to recall the input with a delay of $k$. For both tasks, we use a random binary input $u(t)\in\{0,1\}$.
Figure 5: Information processing capacity $r^{2}$ for the NARMA and delay tasks. (a) $r^{2}$ for the NARMA task as a function of $\sigma_{\text{in}}$ and $\rho$ with $N=500$. (b) $r^{2}$ for the NARMA task with $\rho=0.8$ and various values of $N$. (c) $r^{2}$ for the delay task as a function of $\sigma_{\text{in}}$ and $\rho$ with $N=500$. The value of delay is $k=30$. (d) $r^{2}$ for the delay task with $\rho=0.8$ and various values of $N$. The regions labeled (1) and (2) indicate the linear and nonlinear information processing phases, respectively. Panels (b) and (d) show that the boundary between phases (1) and (2) becomes shaper as $N$ increases.
Figures 5(a) and (b) show the capacity $r^{2}$ for the NARMA task. The blue solid curve in (a) indicates the boundary of the order-to-chaos transition. The regions labeled (1) and (2) correspond to the linear and nonlinear information processing phases, respectively. In contrast to the XOR task, the NARMA task yields a nonzero $r^{2}$ even within the linear regime. Nevertheless, a sharp transition from phase (1) to phase (2) is still clearly observed. Furthermore, as discussed in Appendix C, when $r^{2}$ is plotted against $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$, the curves for different noise levels collapse onto a single curve, confirming the universality of the scaling behavior Eq. (13).
@@
Figures 5(c) and (d) show the capacity $r^{2}$ for the delay task with a delay parameter $k=30$. It is well known that there exists a trade-off between nonlinearity and memory capacity Dambre-12 ; Inubushi-17 , the latter quantifying how well the system retains information about past inputs. Since the delay task relies purely on memory, linear networks tend to achieve higher $r^{2}$ values than their nonlinear counterparts. As a result, $r^{2}$ drops to zero when the system transitions from the linear to the nonlinear regime. In contrast to the XOR and NARMA tasks, we find that the critical value of the unscaled input standard deviation $\sigma_{\text{in}}$ at which the transition occurs is independent of the noise strength $\sigma_{\xi}$ (see Appendix C). This suggests that the transition in the delay task may belong to a different universality class than those observed in the XOR and NARMA tasks.
@@ V Conclusion
In this study, we investigated the transition from linear to nonlinear information processing in neural networks. Our results suggest that the information processing capacity exhibits a discontinuous jump in the limit of infinitely many nodes when plotted against an appropriately rescaled nonlinearity parameter. This transition is distinct from conventional order-chaos bifurcations and equilibrium thermodynamic phase transitions, and thus may enlarge the catalog of nonequilibrium phase transitions in high-dimensional dynamical systems.
@@ Appendix A Training procedure of echo state networks
Figure 7: Information processing capacity $r^{2}$ as a function of $\sigma_{\text{in}}$ for the NARMA task [panels (a) and (b)] and the delay task [panels (c) and (d)]. The number of nodes is $N=600$, and the spectral radius of $W$ is set to $\rho=0.8$. The delay task is performed with a delay of $k=30$. For the NARMA task, the curves corresponding to different noise intensities collapse onto a single master curve when $\sigma_{\text{in}}$ is rescaled as $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. In contrast, for the delay task, a similar collapse is observed without any rescaling of $\sigma_{\text{in}}$.
Appendix B Ridge regression
@@
Appendix C Noise scaling for the NARMA and delay tasks
@@
In this appendix, we discuss how noise influences the information processing capacity $r^{2}$ for the NARMA and delay tasks. Figure 7(a) shows $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for the NARMA task. A similar trend to that observed in the XOR task [see Fig. 3(a)] is evident, with the exception that $r^{2}$ remains nonzero in the linear regime due to the partially linear nature of the NARMA task. In the linear regime, $r^{2}\simeq 0.6$ and is nearly independent of the noise intensity $\sigma_{\xi}$, indicating that linear information processing is robust against noise and that the transition to the nonlinear regime is not driven by an improved signal-to-noise ratio. Figure 7(b) presents $r^{2}$ as a function of the rescaled input standard deviation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. The collapse of the curves corresponding to different noise levels indicates the same scaling behavior observed for the XOR task.
@@
Figure 7(c) shows $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for the delay task. In this case, $r^{2}$ drops to zero as the system transitions from the linear to the nonlinear information processing regime. This behavior arises because nonlinearity degrades the memory capacity of the network, which is essential for solving the delay task. For a relatively large noise level $\sigma_{\xi}=10^{-6}$, we observe a reduction in $r^{2}$ at small values of $\sigma_{\text{in}}$. This reduction is attributed to the detrimental effect of noise on the signal-to-noise ratio when the input amplitude is low. Importantly, the critical value of $\sigma_{\text{in}}$ at which the transition occurs appears to be independent of the noise level. This observation suggests that, in the case of the delay task, the transition is not influenced by the $L^{2}$ regularization typically introduced via ridge regression.
Appendix D Influence of round-off error

### v2

@@ IV.4 Other tasks
To demonstrate the robustness of the transition discussed above, we evaluate the information processing capacity on three additional benchmark problems:
  1. 1.
    Nonlinear autoregressive moving average (NARMA) task. The target is generated by
    $z(t+1)=$ $\ a_{1}z(t)+a_{2}z(t)\sum_{i=0}^{m-1}z(t-i)$
    $+a_{3}u(t-m+1)u(t)+a_{4},$ (21)
    with parameters $m=10$, $a_{1}=0.3$, $a_{2}=0.05$, $a_{3}=0.2$, and $a_{4}=0.05$ Rodan-11 ; Appeltant-11 . The input is a random binary sequence $u(t)\in\{0,1\}$, as in the XOR task.
  2. 2.
    One-step prediction task of the Lorenz system. We consider the Lorenz system, a canonical example of chaotic dynamics, defined by
    $\frac{dX}{dt}$ $=\sigma(Y-X),$
    $\frac{dY}{dt}$ $=X(\rho-Z)-Y,$ (22)
    $\frac{dZ}{dt}$ $=XY-\beta Z,$
    with parameters $\rho=28$, $\sigma=10$, and $\beta=8/3$. The input sequence $u(t)$ is generated by sampling $X(t)$ at intervals of $0.2$, adding Gaussian noise with standard deviation $0.1$ to simulate observation noise, and then normalizing to zero mean and unit variance. The target is set to $z(t)=u(t+1)$, corresponding to a one-step-ahead prediction task based solely on the time series of $X$, rather than the full state $(X,Y,Z)$. This setup reflects realistic scenarios of predicting partially observed chaotic time series.
  3. 3.
    Delay task. The target is defined as z(t)=u(tk)z(t)=u(t-k), which simply requires the network to recall the input with a delay of $k$. We again use a random binary sequence $u(t)\in\{0,1\}$.
Figure 5: Information processing capacity $r^{2}$ for the NARMA, Lorenz, and delay tasks. (a) $r^{2}$ for the NARMA task as a function of $\sigma_{\text{in}}$ and $\rho$ with $N=500$. (b) $r^{2}$ for the NARMA task with $\rho=0.8$ and various values of $N$. (c) $r^{2}$ for the Lorenz task as a function of $\sigma_{\text{in}}$ and $\rho$ with $N=500$. (d) $r^{2}$ for the Lorenz task with $\rho=0.8$ and various values of $N$. (e) $r^{2}$ for the delay task as a function of $\sigma_{\text{in}}$ and $\rho$ with $N=500$. The value of delay is $k=30$. (f) $r^{2}$ for the delay task with $\rho=0.8$ and various values of $N$. The regions labeled (1) and (2) indicate the linear and nonlinear information processing phases, respectively. Panels (b), (d), and (f) show that the boundary between phases (1) and (2) becomes shaper as $N$ increases.
Figures 5(a) and (b) show the capacity $r^{2}$ for the NARMA task. The blue solid curve in (a) indicates the boundary of the order-to-chaos transition. The regions labeled (1) and (2) correspond to the linear and nonlinear information processing phases, respectively. In contrast to the XOR task, the NARMA task yields a nonzero $r^{2}$ even within the linear regime. Nevertheless, a sharp transition from phase (1) to phase (2) is still clearly observed. Furthermore, as discussed in Appendix D, when $r^{2}$ is plotted against $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$, the curves for different noise levels collapse onto a single curve, confirming the universality of the scaling behavior Eq. (13).
@@
Figures 5(c) and (d) show the capacity $r^{2}$ for the one-step prediction task of the Lorenz system. As in the NARMA task, $r^{2}$ remains nonzero even within the linear regime, indicated by label (1). In contrast to the XOR and NARMA tasks, the nonlinear regime [label (2)] exhibits a two-step structure, featuring an intermediate plateau in $r^{2}$. Despite this difference, a clear boundary between the linear and nonlinear regimes can be identified, and it follows the same noise scaling as observed for the XOR and NARMA tasks (see Appendix D). Note that the location of the boundary between regimes (1) and (2) is nearly independent of the spectral radius $\rho$ of the recurrent weight matrix. In general, the spectral radius $\rho$ governs the characteristic time scale of the network dynamics; as $\rho\to 1$, the relaxation time of the undriven network diverges. The nearly vertical phase boundary observed in Fig. 5(c) therefore suggests that the Lorenz prediction task does not require long-term memory.
Figures 5(e) and (f) show the capacity $r^{2}$ for the delay task with a delay parameter $k=30$. It is well known that there exists a trade-off between nonlinearity and memory capacity Dambre-12 ; Inubushi-17 , the latter quantifying how well the system retains information about past inputs. Since the delay task relies purely on memory, linear networks tend to achieve higher $r^{2}$ values than their nonlinear counterparts. As a result, $r^{2}$ drops to zero when the system transitions from the linear to the nonlinear regime. In contrast to the XOR and NARMA tasks, we find that the critical value of the unscaled input standard deviation $\sigma_{\text{in}}$ at which the transition occurs is independent of the noise strength $\sigma_{\xi}$ (see Appendix D). This suggests that the transition in the delay task may belong to a different universality class than those observed in the XOR and NARMA tasks.
@@ V Conclusion
In this study, we investigated the phase transition from linear to nonlinear information processing in neural networks. Our results suggest that the information processing capacity exhibits a discontinuous jump in the limit of infinitely many nodes when plotted against an appropriately rescaled nonlinearity parameter. This transition is distinct from conventional order-chaos bifurcations and equilibrium thermodynamic phase transitions, and thus may enlarge the catalog of nonequilibrium phase transitions in high-dimensional dynamical systems.
At present, we do not have a theoretical explanation for the phase transition reported in this study. A commonly used analytical approach for recurrent neural networks with random weights is the dynamical mean-field theory Sompolinsky-88 ; Kadmon-15 ; Schuecker-18 . For instance, Ref. Schuecker-18 applied this approach to compute the information processing capacity for the delay task in a continuous-time random neural network. However, to the best of our knowledge, no singular behavior of the capacity, as observed in our work, has been reported within this framework. Developing a theoretical understanding of this transition remains an important direction for future research.
In our study, the nonlinearity of the network is controlled by the standard deviation of the input weights. Alternatively, one can consider introducing nonlinearity by incorporating a small number of nonlinear nodes into an otherwise linear network. Our results suggest that even a small fraction of nonlinear nodes can substantially enhance the computational capability of such a network, consistent with the qualitative findings reported in Ref. Inubushi-17 . An intriguing open question is whether there exists a sharp transition at a critical fraction $f_{c}$ of nonlinear nodes, below which the network effectively behaves as linear and above which it achieves an information processing capacity comparable to a fully nonlinear network. Investigating this possibility can provide valuable guidance for identifying the optimal balance between linear and nonlinear nodes in the design of reservoir computing systems.
@@
A low noise level is crucial for observing the transition between linear and nonlinear information processing regimes in physical implementations of reservoir computing. Note that the transition occurs when the rescaled input strength $\tilde{\sigma}_{\text{in},0}=\sigma_{\text{in},0}/\sigma_{\xi}^{1/3}$ is of order unity. Consequently, to achieve a given transition point $\sigma_{\text{in},0}$ requires the noise intensity to scale as $\sigma_{\xi}\sim\sigma_{\text{in},0}^{3}$. In the linear regime, the network state must remain small ($|x_{n}|\ll 1$), so that the activation function is effectively linear ($h(x)\simeq x$). Since the state amplitude is proportional to the input strength $\sigma_{\text{in}}$, the critical input strength must satisfy $\sigma_{\text{in},0}\ll 1$. Consequently, the signal-to-noise ratio required to observe the transition can be estimated as $\sigma_{\text{in},0}/\sigma_{\xi}\sim\sigma_{\text{in},0}^{-2}$, which is typically on the order of $10^{2}$ or higher.
@@ Acknowledgements.
DATA AVAILABILITY
The data that support the findings of this article are openly available data .
@@ Appendix A Training procedure of echo state networks
Appendix B Details of network size scaling
Figure 6: (a) Network size dependence of the maximum derivative $\Delta r^{2}_{\text{max}}$ at the transition point in a double-log plot. The values of the XOR delay are $k=2,4,6,\ldots,24$ from top to bottom. The dashed lines represent $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$. (b) Enlarged view of panel (a), highlighting the power-law dependence. (c) Network size dependence of the transition point $\tilde{\sigma}_{\text{in},0}$ in a double-log plot. The values of the XOR delay are $k=2,4,6,\ldots,24$ from bottom to top. The dashed lines represent $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$. (d) Enlarged view of panel (c). (e), (f) Exponents $\alpha$ and $\beta$ as functions of the delay $k$, obtained using different fitting ranges. The power-law fits for $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ are performed for $N\in[N_{\text{fit,min}},6000]$, with $N_{\text{fit,min}}=2000$, $3000$, and $4000$. The error bars indicate the standard errors from the fitting.
In this appendix, we provide a detailed analysis of how the transition depends on the network size. We focus on two key quantities: the rescaled transition point $\tilde{\sigma}_{\text{in},0}$, defined as the value of $\tilde{\sigma}_{\text{in}}$ at which the derivative $dr^{2}/d\tilde{\sigma}_{\text{in}}$ attains its maximum, and the maximum derivative itself, denoted $\Delta r^{2}_{\text{max}}$. Our aim is to examine how both $\tilde{\sigma}_{\text{in},0}$ and $\Delta r^{2}_{\text{max}}$ scale with the network size $N$.
Figure 6(a) shows $\Delta r^{2}_{\text{max}}$ as a function of $N$ for the various values of the XOR delay $k=2,4,6,\ldots,24$ from top to bottom. The noise intensity is $\sigma_{\xi}=10^{-8}$, and the spectral radius of $W$ is $\rho=0.8$. An enlarged view is provided in Fig. 6(b) to emphasize the power-law scaling behavior. Figure 6(c) shows $\tilde{\sigma}_{\text{in},0}$ as a function of $N$ for the various values of the XOR delay $k=2,4,6,\ldots,24$ from bottom to top. The corresponding magnified view is shown in Fig. 6(d). For sufficiently large $N$, both $\Delta r^{2}_{\text{max}}$ and $\tilde{\sigma}_{\text{in},0}$ exhibit power-law scaling as $\Delta r^{2}_{\text{max}}\propto N^{\alpha}$ and $\tilde{\sigma}_{\text{in},0}\propto N^{-\beta}$ with positive exponents $\alpha$ and $\beta$.
The key questions are: (i) how the scaling exponents $\alpha$ and $\beta$ depend on the XOR delay $k$, and (ii) whether $\alpha$ is larger than, equal to, or smaller than $\beta$. The first question concerns the universality of the scaling exponents. If $\alpha$ and $\beta$ are independent of the XOR delay, this would suggest that they are intrinsic properties of the system, rather than those of the specific benchmark task used to evaluate information processing capacity. The second question addresses the emergence of singular behavior in the limit of infinite system size. If $\alpha=\beta$, then the location and sharpness of the transition scale identically, suggesting a scaling relation of the form $r^{2}(\tilde{\sigma}_{\text{in}},N)=f(\tilde{\sigma}_{\text{in}}N^{\alpha})$ with a smooth function $f(x)$ connecting the linear and nonlinear capacities. By contrast, if $\alpha>\beta$, the transition sharpens more rapidly than the transition point shifts. This would imply that, when plotting $r^{2}$ against $\tilde{\sigma}_{\text{in}}N^{\beta}$, a discontinuous transition may emerge in the limit $N\to\infty$.
Reaching definitive conclusions on the above questions is challenging because the minimal network size $N$ required for $\Delta r^{2}_{\text{max}}$ and $\tilde{\sigma}_{\text{in},0}$ to exhibit clear power-law behavior increases with the XOR delay $k$. As a result, for a fixed fitting range $N\in[N_{\text{fit,min}},N_{\text{fit,max}}]$, the estimated exponents $\alpha$ and $\beta$ may display an artificial $k$-dependence due to transient effects, even if they are in fact constant. Figures 6(e) and (f) show $\alpha$ and $\beta$ as functions of $k$ for fitting ranges $N_{\text{fit,min}},6000]$ with $N_{\text{fit,min}}=2000$, $3000$, and $4000$. A naive interpretation suggests that both $\alpha$ and $\beta$ increase with $k$, and $\alpha>\beta$. However, as $N_{\text{fit,min}}$ increases, the values of $\alpha$ and $\beta$ decrease for large $k$, resulting in a weaker $k$-dependence. If this trend continues at larger $N$, it is possible that both exponents converge to a common universal value of approximately $0.18$, independent of $k$. Given the limitations of numerically accessible system sizes, we cannot rule out this possibility. Further investigations with larger networks will be required to settle this question conclusively.
Appendix C Ridge regression
@@
Appendix D Noise scaling for other tasks
Figure 8: Information processing capacity $r^{2}$ as a function of $\sigma_{\text{in}}$ for the NARMA task [panels (a) and (b)], the Lorenz task [panels (c) and (d)], and the delay task [panels (e) and (f)]. The number of nodes is $N=600$, and the spectral radius of $W$ is set to $\rho=0.8$. The delay task is performed with a delay of $k=30$. For the NARMA and Lorenz tasks, the curves corresponding to different noise intensities collapse onto a single master curve when $\sigma_{\text{in}}$ is rescaled as $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. In contrast, for the delay task, a similar collapse is observed without any rescaling of $\sigma_{\text{in}}$.
@@ Appendix D Noise scaling for other tasks
In this appendix, we discuss how noise influences the information processing capacity $r^{2}$ for the NARMA, Lorenz, and delay tasks. Figure 8(a) shows $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for the NARMA task. A similar trend to that observed in the XOR task [see Fig. 3(a)] is evident, with the exception that $r^{2}$ remains nonzero in the linear regime due to the partially linear nature of the NARMA task. In the linear regime, $r^{2}\simeq 0.6$ and is nearly independent of the noise intensity $\sigma_{\xi}$, indicating that linear information processing is robust against noise and that the transition to the nonlinear regime is not driven by an improved signal-to-noise ratio. Figure 8(b) presents $r^{2}$ as a function of the rescaled input standard deviation $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. The collapse of the curves corresponding to different noise levels indicates the same scaling behavior observed for the XOR task. The same behavior can also be observed for the Lorenz task, as shown in Figs. 8(c) and (d).
@@
Figure 8(e) shows $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for the delay task. In this case, $r^{2}$ drops to zero as the system transitions from the linear to the nonlinear information processing regime. This behavior arises because nonlinearity degrades the memory capacity of the network, which is essential for solving the delay task. For a relatively large noise level $\sigma_{\xi}=10^{-6}$, we observe a reduction in $r^{2}$ at small values of $\sigma_{\text{in}}$. This reduction is attributed to the detrimental effect of noise on the signal-to-noise ratio when the input amplitude is low. Importantly, the critical value of $\sigma_{\text{in}}$ at which the transition occurs appears to be independent of the noise level. This observation suggests that, in the case of the delay task, the transition is not influenced by the $L^{2}$ regularization typically introduced via ridge regression.
Appendix E Influence of activation function
Figure 9: Information processing capacity $r^{2}$ as a function of $\sigma_{\text{in}}$ for the XOR task with different activation functions $h(x)$. Panels (a) and (b) correspond to the error function $h(x)=\text{erf}(\sqrt{\pi}x/2)$, (c) and (d) to a fifth-order function $h(x)=x+O(x^{5})$, and (e) to a piecewise linear function. The corresponding activation functions are plotted in panel (f). The number of nodes is $N=600$, and the spectral radius of $W$ is set to $\rho=0.8$. The delay in the XOR task is $k=10$ for panels (a) and (b), and $k=5$ for panels (c)-(e). For the error function [panels (a) and (b)], the curves for different noise intensities $\sigma_{\xi}$ collapse onto a single master curve when $\sigma_{\text{in}}$ is rescaled as $\sigma_{\text{in}}/\sigma_{\xi}^{1/3}$. For the fifth-order function [panels (c) and (d)], the appropriate rescaling becomes $\sigma_{\text{in}}/\sigma_{\xi}^{1/5}$, reflecting the order of the nonlinearity in $h(x)$. For the piecewise linear function [panel (e)], the capacity $r^{2}$ remains essentially independent of the noise intensity when it is sufficiently small.
In this appendix, we examine how the choice of activation function $h(x)$ in Eq. (6) influences the noise scaling behavior associated with the transition from linear to nonlinear information processing. We consider the following three types of activation functions:
  1. 1.
    Error function.
    $h(x)=\text{erf}\left(\frac{\sqrt{\pi}}{2}x\right),\quad\text{erf}(x):=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt.$ (25)
  2. 2.
    Fifth-order function.
    $h(x)=\begin{cases}-1&(x<-5/4),\\ x-\frac{1}{5}\left(\frac{4}{5}\right)^{4}x^{5}&(-5/4\leq x<5/4),\\ 1&(x\geq 5/4).\end{cases}$ (26)
  3. 3.
    Piecewise linear function.
    $h(x)=\begin{cases}-1&(x<-1),\\ x&(-1\leq x<1),\\ 1&(x\geq 1).\end{cases}$ (27)
Each activation function admits an expansion of the form $h(x)=x+O(x^{\zeta})$ around $x=0$, where $\zeta$ denotes the order of the leading nonlinear correction. Specifically, we have $\zeta=3$ for the error function, $\zeta=5$ for the fifth-order function, and $\zeta=\infty$ for the piecewise linear function. The activation functions are shown in Fig. 9(f). In the remainder of this section, we investigate the relationship between $\zeta$ and the scaling exponent $\eta$, which characterizes how the transition point $\sigma_{\text{in},0}$ scales with the noise intensity $\sigma_{\xi}$.
Figure 9(a) shows the information processing capacity $r^{2}$ as a function of the input standard deviation $\sigma_{\text{in}}$ for the error function activation. As shown in Fig. 9(b), the rescaled curves exhibit the same scaling behavior as those obtained with the $\tanh$ activation, characterized by a scaling exponent $\eta=1/3$. Figures 9(c) and (d) show $r^{2}$ as a function of $\sigma_{\text{in}}$ and its rescaled counterpart, respectively, for the fifth-order activation function. In this case, the observed scaling exponent is $\eta=1/5$, reflecting the higher-order nonlinearity of the activation function. Figure 9(e) shows $r^{2}$ for the piecewise linear activation function, where the capacity remains independent of $\sigma_{\xi}$ when it is sufficiently small, indicating that $\eta=0$. These results support the hypothesis that the scaling exponent $\eta$ is determined by the inverse of the leading-order nonlinearity $\zeta$ in the expansion of the activation function, i.e., $\eta=1/\zeta$.
Appendix F Influence of round-off error